OpenStudy (anonymous):

5/x+4 + 1/x+4= x-1 (solve for x) HELP!

OpenStudy (anonymous):

\[\frac{5}{x+4}+\frac{1}{x+4}=x-1\] like that?

OpenStudy (anonymous):

add to get \[\frac{6}{x+4}=x-1\] so \[6=(x+4)(x-1)\] multiply out, get a quadratic equation in x

OpenStudy (anonymous):

\[x^2+3x-4=6\] \[x^2+3x-10=0\] \[(x+5)(x-2)=0\] \[x=-5,x=2\]